OpenStudy (3psilon):

f(2) = -8 , f'(2)=3 , g(2) = 17 and g'(2) = -4 what is the value of (f g)' (2)

OpenStudy (anonymous):

\[f(g(2))\]

OpenStudy (anonymous):

\[f'(g(2))\] \[f(g'(2))\]

OpenStudy (anonymous):

We can use the chain-rule to figure such out: We know: \[ (f(g(x)))'=f'(g(x))g'(x) \]So, try that, see if helps

OpenStudy (anonymous):

which one is it?

OpenStudy (anonymous):

is that asking the derivative of the product of the two functions?

OpenStudy (anonymous):

That's asking for composition, I think, since we have\[ (f\circ g)'(x) \]

OpenStudy (anonymous):

if it is the composition, then your stuck because the problem does not say anything about \(\large f'(17)= \) ?????

OpenStudy (3psilon):

@LolWolf what would f'(17) be?

OpenStudy (3psilon):

-204?

OpenStudy (anonymous):

I think it's supposed to be a product? (f*g) ' ?

OpenStudy (3psilon):

I did and thats what I got @Algebraic!

OpenStudy (anonymous):

I think earlier replies were assuming it was a composition...

OpenStudy (3psilon):

What is a composition ?

OpenStudy (anonymous):

f( g(x) )

OpenStudy (anonymous):

as opposed to f(x) * g(x)

OpenStudy (anonymous):

answer is 83?

OpenStudy (3psilon):

Can you show me how you got there?

OpenStudy (anonymous):

product rule : (f*g) ' = f' *g + f* g'

OpenStudy (anonymous):

=3*17 + (-8)*(-4)

OpenStudy (3psilon):

And it is (fg)' (2) so 166?

OpenStudy (anonymous):

83

OpenStudy (3psilon):

But the question is Value of (fg)' (2) what do you do with the 2 then?

OpenStudy (anonymous):

naw, that's the value of x at which they want you to evaluate the expression

OpenStudy (anonymous):

ie at x=2 (f*g) ' = ...?

OpenStudy (anonymous):

as far as I can tell from your problem statement...

OpenStudy (3psilon):

Okay I see what you did there! Thanks

OpenStudy (anonymous):

np:)